Francis, LLC is considering a 4-year project with an initial cost of $120,000. The equipment depreciation is straight-line to zero over the life of the project. The firm believes that they can sell the equipment for the salvage value of $40,000 at the end of the project. The tax rate is 40%. The project requires no additional working capital over the life of the project. Annual operating cash flows are $48,000. Assuming a required rate of return of 10%, what is the NPV of this project?
Answer #1

Annual depreciation = 120,000 / 4 = 30,000

Year 4 non operating cash = Market value - tax(market value - book value)

Year 4 non operating cash = 40,000 - 0.4(40,000 - 0)

Year 4 non operating cash = 40,000 - 16,000

Year 4 non operating cash = 24,000

NPV = Present value of cash inflows - present value of cash outflows

NPV = -120,000 + 48,000 / (1 + 0.1)1 + 48,000 / (1 + 0.1)2 + 48,000 / (1 + 0.1)3 + 48,000 / (1 + 0.1)4 + 24,000 / (1 + 0.1)4

NPV = $48,545.86